OpenStudy (anonymous):

What is the simplified form of the reciprocal of negative squaroot 48?

OpenStudy (anonymous):

\[-\frac{1}{\sqrt{48}}\] \[-\frac{1}{\sqrt{16\times 3}}\] \[-\frac{1}{4\sqrt{3}}\] \[-\frac{1}{4\sqrt{3}}\times \frac{\sqrt{3}}{\sqrt{3}}\]

OpenStudy (anonymous):

\[-\frac{\sqrt{3}}{12}\]

OpenStudy (anonymous):

ur smart... i feel so dumb

OpenStudy (anonymous):

you get used to doing these. multiply top and bottom by radical to get it out of denominator

OpenStudy (anonymous):

Can you help me with this problem.... Simplify: squareroot of 0.6

OpenStudy (anonymous):

always look for a perfect square inside radial. \[\sqrt{50}=\sqrt{25\times 2}=5\sqrt{2}\]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

\[\sqrt{.6}=\sqrt{\frac{6}{10}}=\sqrt{\frac{3}{5}}\]

OpenStudy (anonymous):

now rationalize the denominator \[\frac{\sqrt{3}}{\sqrt{5}}\times\frac{\sqrt{5}}{\sqrt{5}}=\frac{\sqrt{15}}{5}\]\]

OpenStudy (anonymous):

do you know how to solve this one? If 3x + 4y = 7 and 2x-3y=67 then which of the following is the value of xy?

OpenStudy (anonymous):

sure you have to solve the system and then multiply

OpenStudy (anonymous):

can you solve this system?

OpenStudy (anonymous):

wait what do you mean by solving the system and multiplying